OpenStudy (anonymous):

Find the perimeter of each of the two noncongruent triangles where a=15 b=20 A=29

OpenStudy (ddcamp):

Law of sines to find possible B values: \[\frac{ Sin(A) }{ a } = \frac{ Sin(B) }{ b }\] Then, using these two B values you can find the two possible C values \[A° + B° + C° = 180°\] And the law of cosines to find the length of c: \[c^2 = a^2 + b^2 -2(ab)\cos(C)\]